# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 2.1 Definitions
  • We write $\mathcal{B}\langle X,X^{\ast}\rangle$ for the algebra generated by $\mathcal{B}$ and the freely non-commuting indeterminates $(X,X^{\ast})$. For a monomial $P(X,X^{\ast})=b_{0}X^{\epsilon_{1}}b_{1}X^{\epsilon_{2}}\cdots X^{\epsilon_{n}}b_{n}\in\mathcal{B}\langle X,X^{\ast}\rangle$ the degree of $P(X,X^{\ast})$ is $n$, its coefficients are $b_{0},...,b_{n}\in\mathcal{B}$ and for all $1\leq k\leq n$ we have $\epsilon_{k}\in\{+,-\}$. Then the operator valued distribution, or $\mathrm{E}$-distribution, of $a\in\mathcal{A}$ is the map given by:
    $\mathrm{E}_{a}\leavevmode\nobreak\ :\leavevmode\nobreak\ P(\cdot)\in\mathcal{B}\langle X,X^{\ast}\rangle\mapsto\mathrm{E}[P(a,a^{\ast})]\in\mathcal{B}.$
@@
  • The triple $(\mathcal{A},\Phi,\ast)$ is a non-commutative probability space with a faithfull trace, i.e. we have $\Phi:\mathcal{A}\rightarrow\operatorname{\mathbb{C}}$ linear, verifying $\Phi(1)=1$ and
@@
Note that $(\mathrm{M}_{N}(\mathbb{C}),\mathrm{D}_{N}(\mathbb{C}),\Delta,\operatorname{Tr}_{N},\ast)$ and $(\mathrm{M}_{N}(L^{\infty}),\mathrm{D}_{N}(L^{\infty}),\Delta,\mathbb{E}\circ\operatorname{Tr}_{N},\ast)$ are operator-valued probability spaces. With no loss of generality here, any operator-valued probability space can be viewed as a subalgebra of a $\mathcal{B}(\mathrm{H})$, set of bounded operators on a Hilbert space $\mathrm{H}$. The main result of the present article consists in upper bounding $\rho(A_{\#}|_{\boldsymbol{1}^{\perp}})$ for #$\#\in\{\emptyset,r\}$ defined by (3) and (4) by an approximation of the spectral radius of respectively $a$ and $p\in(\mathcal{A},\mathcal{B},\mathrm{E},\Phi)$, defined by (10) and where $q$ has same $\ast$-distribution as $Q$ and $(q,u)$ are free with amalgamation over $\mathcal{B}$. More precisely, owing to Gelfand’s Theorem we have for such $\mathrm{b}\in\mathcal{B}(\mathrm{H})$
$\rho(\mathrm{b})=\lim_{n\to\infty}\|\mathrm{b}^{n}\|^{1/n}.$
Therefore the approximations of spectral radius we will use are
$\mathrm{N}_{n}(Q)=\|a^{n}\|^{1/n}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{N}_{n,r}(Q)=\|p^{n}\|^{1/n}$ (11)
@@
for a fixed $n\in\operatorname{\mathbb{N}}$. We will omit the dependence on $Q$ and simply denote by $\mathrm{N}_{n}$ or $\mathrm{N}_{r,n}$ when there is no confusion. The fact that this norm depends only on the matrix $Q$ will be a consequence of the construction of the operators $(q,u)$ in Section 4.
@@ 2.2 Main result
We fix $10<r<$ 1 and we set
$\rho_{n}=\max\{\mathrm{N}_{n};2\kappa(Q);\delta\}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho_{n,r}=\max\{\mathrm{N}_{n,r};2r\kappa(Q);r\delta\}$ (12)
where $\kappa(Q)$ is given by (2) and $\mathrm{N}_{n,\#}$ by (11) for #$\#\in\{\emptyset,r\}$. The main theorem is the following one.
@@
  • $q$ and $Q$ have same $\ast$-distribution,
  • $q$ and $u$ are free with amalgamation over $\mathcal{B}$.
We further assume that $Q$ verifies (H:1), (H:2) and (H:3). Then for any $n$ integer and $10<c<$ 1 there exists $C>0$ such that setting
@@
$\mathbb{P}(\rho(A_{\#}|_{\boldsymbol{1}^{\perp}})\geq(1+\epsilon)\rho_{n,\#})\leq N^{-c},$
for $M\sim\mathcal{U}(\mathrm{S}_{N})$, #$\#\in\{\emptyset,r\}$ and $N$ large enough.
Acknowledgments
@@ 3 Some applications
In this section we give examples where 2.2 holds, and where the upper bound $\rho_{n,\#}$ defined by (12) is equal to the approximation of the spectral radius of the deterministic operator. We also give an example where an improvement of $\rho_{n}$ would be appreciated to conclude.
@@ 3.1 Convex combination
In the setting of 2.2 and for any $Q\in\mathrm{M}_{N}(\mathbb{C})$ that verifies (H:1), it is sufficient to consider $10<r<$ 1 small enough to have $\rho_{n,r}=\|[(1-r)u+rq]^{n}\|^{1/n}$ where $\rho_{n}$ is defined by (12). Indeed we have then $Q_{xy}\leq\delta$ for all $x,y\in[N]$ and therefore for any integers $1\leq k\leq p$ and $(n_{t})_{t\leq k}$ such that $\sum_{t=1}^{k}n_{t}=p$ and for any $(x_{t},y_{t})_{t}\in[N]^{2k}$ we have
@@
which implies $\kappa(Q)\leq(1\vee\delta)$ where $\kappa(\cdot)$ is defined by (2). Furthermore owing to 4.7, we have that $\|[(1-r)u+rq]^{n}\|^{1/n}\geq(1-r)$. We deduce that for all $0<r<\frac{1}{3(\delta\vee 1)}$ we have $\rho_{n}=\|[(1-r)u+rq]^{n}\|^{1/n}$ as desired. In particular for $Q$ bistochastic the upper bound $\rho_{n}$ is the approximation $\rho_{n}=\|[(1-r)u+rq]^{n}\|^{1/n}$ for any $n\geq 0$ and for any $0<r<\frac{1}{3}$. To then show that the spectral gap is strictly positive, one would then need to provide an upper bound for the norm $\|[(1-r)u+rq]^{n}\|^{1/n}$ that is strictly less than $1$. In the following section we give an explicit example where we can give such an upper bound.
@@ 3.2 The asymptotically free case
We consider $Q$ the Markov operator of a simple walk obtain with a perfect matching. One way to compute $Q$ is to take any $S\in\mathrm{S}_{N}$ permutation matrix and consider $Q=SUS^{\ast}$ where $U=\operatorname{Diag}\left(\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\cdots\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right)$. The spectral gap for the simple walk defined by $Q$ is trivially equal to 00 since $1$ is an eigenvalue with an eigenspace of dimension at least $2$ for $N\geq 4$. Computing the result of Biane and Lehner [BL99] we obtain that the spectral radius of $a=u+q$ where $q$ has same distribution as $Q$ and $u$ is free with $Q$ is $\rho=\sqrt{3}$ (given by the equation $|\lambda|^{2}+1=|\lambda^{2}-1|^{2}$). For any $\epsilon>0$, owing to Gelfand theorem, there exists $n_{0}>0$ such that for all $n\geq n_{0}$ we have $\mathrm{N}_{n}\leq\left(1+\epsilon\right)\rho$. A realization of the spectrum of $A=M+Q$ encircled by the circle of radius $\rho$ is given in Figure 1.
@@
However the 2.2 does not allow to directly conclude about the spectral gap for $A$. Indeed in this case we have $\kappa=1$ and therefore $\rho_{n}=2\kappa=\lambda_{1}(A)$. It shows that an improvement on the factor $2$ of $\kappa$ in the definition of $\rho_{n}$ would be beneficial to our purposes. This factor appears in the proof of 7.5, Equation 33.
However, one can consider again the convex combination case (i.e. the matrix $A_{r}$) one can compute the explicit asymptotical radius applying Biane and Lehner method. We rewrite
$p=(1-r)u+rq=(1-r)\left(u+\frac{r}{1-r}q\right)=(1-r)(u+q^{\prime})$
@@
where $\varphi=r/(1-r)$ (see [BL99, Equation (4.5)]). In this case, owing to 4.7 we have $\mathrm{N}_{r,n}\geq(1-r)$ for all $n$ and therefore considering $r\leq 1/3$ we have $\rho_{n,r}=\mathrm{N}_{n,r}$. Besides, for all $\epsilon>0$ we have for $n$ large enough $N_{n,r}\leq\rho(p)+\epsilon$ which proves a strictly positive lower bound for the spectral gap for $N$ large enough.
@@
Figure 2: In blue the eigenvalues of $P=(1-r)M+rQ$ with $Q$ perfect matching and $N=600$. In red the spectral radius of $p=(1-r)u+rq$.

### v4

@@ 1.4 Free probability and spectral analysis of random matrices
  • We write $\mathcal{B}\langle X,X^{\ast}\rangle$ for the algebra generated by $\mathcal{B}$ and the freely non-commuting indeterminates $(X,X^{\ast})$. For a monomial $P(X,X^{\ast})=b_{0}X^{\epsilon_{1}}b_{1}X^{\epsilon_{2}}\cdots X^{\epsilon_{n}}b_{n}\in\mathcal{B}\langle X,X^{\ast}\rangle$ the degree of $P(X,X^{\ast})$ is $n$, its coefficients are $b_{0},...,b_{n}\in\mathcal{B}$ and for all $1\leq k\leq n$ we have $\epsilon_{k}\in\{+,-\}$. Then the operator valued distribution, or $\mathrm{E}$-distribution, of $a\in\mathcal{A}$ is the map given by
    $\mathrm{E}_{a}\penalty 10000\ :\penalty 10000\ P(\cdot)\in\mathcal{B}\langle X,X^{\ast}\rangle\mapsto\mathrm{E}[P(a,a^{\ast})]\in\mathcal{B}.$
@@
  • The triple $(\mathcal{A},\Phi,\ast)$ is a non-commutative probability space with a faithful trace, i.e. we have $\Phi:\mathcal{A}\rightarrow\operatorname{\mathbb{C}}$ linear, verifying $\Phi(1)=1$ and
@@
Note that $(\mathrm{M}_{N}(\mathbb{C}),\mathrm{D}_{N}(\mathbb{C}),\Delta,\operatorname{Tr}_{N},\ast)$ and $(\mathrm{M}_{N}(L^{\infty}),\mathrm{D}_{N}(L^{\infty}),\Delta,\mathbb{E}\circ\operatorname{Tr}_{N},\ast)$ are operator-valued probability spaces. With no loss of generality here, any operator-valued probability space can be viewed as a subalgebra of a $\mathrm{B}(\mathrm{H})$, set of bounded operators on a Hilbert space $\mathrm{H}$. This freeness has a particular interest in random graph theory because the previous theorem holds again replacing freeness with freeness with amalgamation and invariance under $\mathrm{G}_{N}$ conjugation by invariance under $\mathrm{S}_{N}$ conjugation.
Theorem 1.3.
[ACD+18, Theorem 1.3] If $Q\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ converges in distribution and is invariant under $\mathrm{S}_{N}$ conjugation and $M\in\mathrm{S}_{N}$ is uniformly distributed, then the couple $(Q,M)$ is asymptotically free with amalgamation over the diagonal, i.e the subalgebra of diagonal matrices.
Once again the convergence in distribution and the freeness with amalgamation of the limit operators allows to compute the distribution of the limit family. An explicit algorithm to compute the joint distribution of free with amalgamation operators is given by Belinschi, Mai, Speicher [BMS17]. The previous setting is better suited for graph theory applications, since we can compute random graphs which joint entry wise law is not invariant under $\mathrm{O}_{N}$ or $\mathrm{U}_{N}$ conjugation but under $\mathrm{S}_{N}$. Indeed one can consider $Q$ the adjacency matrix (resp. the Markov operator of the simple walk) of an Erdős-Rényi graph, then the couple $(Q,M)$ is asymptotically free with amalgamation and the asymptotic spectral distribution of the operators defined by (4) and (5) can be computed. Again, in an Hermitian case where we consider $Q$ a deterministic hermitian matrix invariant under $\mathrm{S}_{N}$ conjugation and $M$ a uniformly distributed random permutation, considering
$\mathrm{P}:=Q+M+M^{\ast}$
one can apply ˜1.3 and denoting $(q,u)$ the limit free with amalgamation operators for the joint distribution of $(Q,M)$ we have again
$|\sigma(\mathrm{P})\cap[-\|\mathrm{b}\|,\|\mathrm{b}\|]^{\mathrm{c}}|=o(N)$
where $\mathrm{b}=q+u+u^{\ast}$.
In our case, not only we do not assume any asymptotic freeness of the couple $(Q,M)$, but we also consider non-Hermitian operators (4) (5). Nevertheless we show that we have expansion under some assumptions on $Q$ distribution ((H:1), (H:2), (H:3)). The fact that we use an approximation of the spectral radius of the sum of free with amalgamation operators will then allow us to say that the expansion we prove here is optimal, up to showing the weak convergence of the empirical spectral measure towards the brown measure of the limit operator. Computing the brown measure radius for the limit object is possible in some cases [BL99]. Also simulations lead to think that it is indeed this radius that gives the asymptotic second largest eigenvalue in modulus (see Section˜4, Figure˜1 and Figure˜2).
Whether one has asymptotic freeness or freeness with amalgamation of $(Q,M)$, the fact that we are dealing with non-Hermitian operators invalidates the spectral theorem for the operator p=(1r)u+rqp=(1-r)u+rq. One can then ask if Girko’s method applies to show the convergence (7) with $\mu_{\mathrm{b}}$ the brown measure of the operator $\mathrm{b}$ (see for instance [BC12, RV08] for $Q$ with independent entries, [BCZ18] for $P=\frac{1}{d}\sum_{s}M(s)$ with $(\ln(N)^{12}/\ln(\ln(N))^{4}\leq d=o(N)$). This convergence would then depends on the distribution chosen for $Q$. In the case where we have the convergence (7) we then can conclude that the expansion given in this article is optimal in regard of (8).
Strong asymptotic freeness.
In any case, the convergence given in (7) is not sufficient to conclude about the expansion of operators. In fact, it is precisely the absence of outliers that is needed to satisfy the expander ˜1.1. One way to obtain these expansions is to show strong asymptotic freeness. This strong convergence had been shown for $\mathbf{M}=(M(s))_{s\in[d]}$ a family of iid uniformly distributed random permutation by Bordenave and Collins [BC19]. In particular it means that the random regular graph defined by the adjacency operator $A=\sum_{s=1}^{d}M(s)+M(s)^{\ast}$ has the best expansion possible, i.e it verifies that for all $\epsilon>0$ we have that for $N$ large enough $\lambda_{2}(A)\leq 2\sqrt{2d-1}+\epsilon$. The fact that the best expansion possible is given by $2\sqrt{2d-1}$ can be seen as a consequence of the asymptotic freeness given by Nica [NIC93] as the spectral radius of the limit operator $\mathrm{b}$ in the convergence (7), or directly by the Alon-Boppana bound of regular graph (see for instance [MOH10]). We now give a simplified version of ˜2.1.
Theorem 1.4.
Let $Q$ be a bistochastic matrix and $M$ an independent uniformly distributed random permutation. There exists $(q,u)$ a pair of free with amalgamation operator such that $u$ is unitary and $q$ and $Q$ have same $\ast$-distribution asymptotically. Besides for any $0<r<1/3$ and $0<c<1$ and for all $\epsilon>0$ we have
$\mathbb{P}(\left|\lambda_{2}((1-r)M+rQ)\right|\leq(1+\epsilon)\rho((1-r)u+rq))\geq N^{-c}$
for $N$ large enough.
The previous simplified version is the case discussed in Section˜4.1. If one can compute $\rho=\rho((1-r)u+rq)$ in the previous theorem and if $\rho<1$ this gives us expansion for the random walk defined by (1r)M+rQ(1-r)M+rQ. To prove then that this expansion is optimal, one way would be to prove the convergence of the empirical spectral measure towards a measure which support is countained in the complex disc of radius $\rho$. In regards of [BC12, Lemma 4.3] in the case where $(M,Q)$ is asymptotically free with amalgamation, the only thing left to show that the empirical spectral converges towards the brown measure of the operator (1r)u+rq(1-r)u+rq is the $\log$-integrability, which often comes down to controlling singular values of $Q+M$. Finally, an interest of looking directly at the operator $A=Q+M$ is that its structure and the computations of its moments are closer to the more general case where we study $A^{\prime}=a_{0}\otimes\operatorname{Id}_{N}+a_{1}\otimes Q+a_{2}\otimes M$ with $a_{i}$ fixed matrices of dimension $k$. Indeed, in many cases the strong convergence in the free case is showed using a linearization trick saying that if the spectrum of $A^{\prime}$ for any choices of $(a_{i})_{i}$ such that $A^{\prime}$ is Hermitian converges then we also have convergence of the spectrum of $P(M,Q)$ for any symmetric polynomial $P(\cdot)$ (see for instance [BC19, Section 6]). The main issue in adapting this linearization trick to our case is the fact that the matrix $a_{1}\otimes Q$ does not verify (H:1).
2 Main result
2.1 Existence of the free with amalgamation operators
We need the following lemma to enunciate our main theorem.
Lemma 2.1.
  • asymptotically, the operators q{q} and $Q$ have the same $\ast$-distribution,
  • The operators $q$ and $u$ are free with amalgamation over $\mathcal{B}$.
The proof of the previous lemma and useful characteristics of of the couple $(q,u)$ are given in Section˜6. For any matrix $Q\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ we define the operators $\{\mathrm{b}_{\mathrm{x}}\}_{\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}}$ as follows
$\mathrm{b}_{\mathrm{s}}:=u+q,\penalty 10000\ \penalty 10000\ \mathrm{b}_{\mathrm{c}}:=(1-r)u+rq$ (9)
@@ 2.1 Existence of the free with amalgamation operators
where $(q,u)$ are the operators given by ˜2.1. The previous operators asymptotically bounds the spectral radius of the matrices defined by (4) and (5). The main result of the present article consists in upper bounding $\rho(A_{\mathrm{x}}|_{\boldsymbol{1}^{\perp}})$ for $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$ defined by (4) and (5) by an approximation of the spectral radius of $\mathrm{b}_{\mathrm{x}}$ defined by (9). More precisely, owing to Gelfand’s Theorem we have for such $\mathrm{b}\in\mathrm{B}(\mathrm{H})$
$\rho(\mathrm{b})=\lim_{n\to\infty}\|\mathrm{b}^{n}\|^{1/n}$
where one has to recall here that with no loss of generality we can assume that the algebra $\mathcal{A}$ given by ˜2.1 verifies $\mathcal{A}=\mathrm{B}(\mathrm{H})$ for some Hilbert space $\mathrm{H}$. Therefore the approximations of spectral radius we will use are given by
$\mathrm{N}_{n,\mathrm{x}}(Q)=\|\mathrm{b}_{\mathrm{x}}^{n}\|^{1/n}$ (10)
@@
for $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$ and a fixed $n\in\operatorname{\mathbb{N}}$. We will omit the dependence on $Q$ and simply denote by $\mathrm{N}_{n,\mathrm{x}}$ when there is no confusion. The fact that this norm depends only on the matrix $Q$ will be a consequence of the construction of the operators $(q,u)$ in Section˜6.
@@ 2.2 Main result
We fix $0<r<1$ and we set
$\rho_{n,\mathrm{s}}=\max\{\mathrm{N}_{n,\mathrm{s}},2\kappa(Q),\delta,2\|Q_{\perp}\|\}\penalty 10000\ \text{and}\penalty 10000\ \rho_{n,\mathrm{c}}=\max\{\mathrm{N}_{n,\mathrm{c}},2r\kappa(Q),r\delta,2r\|Q_{\perp}\|\}$ (11)
where $\kappa(Q)$ is given by (2) and $\mathrm{N}_{n,\mathrm{x}}$ by (10) for $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$. The main theorem is the following one.
@@
We consider $Q$ a matrix of dimension NN that verifies (H:1), (H:2) and (H:3). Then for any $n$ integer and $0<c<1$ there exists $C>0$ such that setting
@@
$\mathbb{P}(\left|\lambda_{2}(A_{\mathrm{x}})\right|\geq(1+\epsilon)\rho_{n,\mathrm{x}})\leq N^{-c},$
for $M\sim\mathcal{U}(\mathrm{S}_{N})$, $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$ and $N$ large enough.
3 Needed improvements
The strategy of the proof of the main technical statement of this paper, given by ˜7.1 is largely inspired by previous bounding of norms and/or spectral radius using tangle-freeness of paths in the moment method [BOR15, BC19, BDH22]. A good improvement of ˜2.1 would be to get rid of $\delta=\sum_{y}Q_{x,y}$ in the upper bound given by ˜8.4, which would then give the following theorem.
Conjecture 3.1.
˜2.1 holds if one replaces $\rho_{n,\mathrm{x}}$ with
$\tilde{\rho}_{n,\mathrm{s}}=\max\{\mathrm{N}_{n,\mathrm{s}},2\kappa(Q),2\|Q_{\perp}\|\}\penalty 10000\ \text{and}\penalty 10000\ \tilde{\rho}_{n,\mathrm{c}}=\max\{\mathrm{N}_{n,\mathrm{c}},2r\kappa(Q),2r\|Q_{\perp}\|\}.$
An important application of the previous statement would be for $A$ an adjacency matrix of a $d$-regular independent of $M$ a uniformly distributed random matrix. Then we would have that for $N$ large enough, $|\lambda_{2}(A+M)|\leq C\sqrt{d}$ for some universal constant $C$ and therefore we have a good expansion. We give more detail about a possible strategy to obtain this conjecture in ˜8.2.
@@ 4 Some applications
In this section we give examples where ˜2.1 holds, and where the upper bound $\rho_{n,\mathrm{x}}$ defined by (11) is equal to the approximation of the spectral radius of the deterministic operator, i.e ${\mathrm{N}}_{n,\mathrm{x}}$ defined by (10). For these applications it is important to notice that owing to ˜6.4 we have
$\mathrm{{N}}_{n,\mathrm{x}}\geq\mathds{1}(\mathrm{x}=\mathrm{s})+(1-r)\mathds{1}(\mathrm{x}=\mathrm{c}).$
We also give an example where an improvement of $\rho_{n,\mathrm{s}}$ would be appreciated to conclude.
@@ 4.1 Convex combination
In the setting of ˜2.1 and for any $Q\in\mathrm{M}_{N}(\mathbb{C})$ that verifies (H:1), it is sufficient to consider $0<r<1$ small enough to have $\rho_{n,\mathrm{c}}={\mathrm{N}}_{n,\mathrm{c}}$ where $\rho_{n,\mathrm{c}}$ is defined by (11). Indeed we have then $Q_{xy}\leq\delta$ for all $x,y\in[N]$ and therefore for any integers $1\leq k\leq p$ and $(n_{t})_{t\leq k}$ such that $\sum_{t=1}^{k}n_{t}=p$ and for any $(x_{t},y_{t})_{t}\in[N]^{2k}$ we have
@@
which implies $\kappa(Q)\vee\|Q_{\perp}\|\leq 1\vee\delta$ where $\kappa(\cdot)$ is defined by (2). In particular for $Q$ bistochastic the upper bound is the approximation $\rho_{n,\mathrm{c}}=\mathrm{N}_{n,\mathrm{c}}$ for any $n\geq 0$ and for any $0<r<\frac{1}{3}$. To then show that the spectral gap is strictly positive, one would then need to provide an upper bound for the norm $\|[(1-r)u+rq]^{n}\|^{1/n}$ that is strictly less than $1$. In the following application we give an explicit example where we can give such an upper bound. In the formulation of ˜1.4 we used the fact that owing to Gelfand theorem we have that for all $\epsilon>0$ we have $\|[(1-r)u+rq]^{n}\|^{1/n}\geq(1-\epsilon)\rho((1-r)u+rq)$ for $n$ large enough.
@@ 4.2 The asymptotically free case
We consider $Q$ the Markov operator of a simple walk obtain with a perfect matching. One way to compute $Q$ is to take any $S\in\mathrm{S}_{N}$ permutation matrix and consider $Q=SUS^{\ast}$ where $U=\operatorname{Diag}\left(\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\cdots\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right)$. The spectral gap for the simple walk defined by $Q$ is trivially equal to 0 since $1$ is an eigenvalue with an eigenspace of dimension at least $2$ for $N\geq 4$. Computing the result of Biane and Lehner [BL99] we obtain that the spectral radius of $a=u+q$ where $q$ has same distribution as $Q$ and $u$ is free with $Q$ is $\rho=\sqrt{3}$ (given by the equation $|\lambda|^{2}+1=|\lambda^{2}-1|^{2}$). For any $\epsilon>0$, owing to Gelfand theorem, there exists $n_{0}>0$ such that for all $n\geq n_{0}$ we have ${\mathrm{N}}_{n,\mathrm{s}}\leq\left(1+\epsilon\right)\rho$. A realization of the spectrum of $A=M+Q$ encircled by the circle of radius $\rho$ is given in Figure˜1.
@@
However the ˜2.1 does not allow to directly conclude about the spectral gap for $A$. Indeed in this case we have $\kappa=1$ and therefore $\rho_{n,\mathrm{s}}=2\kappa=2\|Q_{\perp}\|=\lambda_{1}(A)$. It shows that an improvement on the factor $2$ of $\kappa$ in the definition of $\rho_{n}$ would be beneficial to our purposes. This factor appears in (38) in the proof of ˜8.4.
However, one can consider again the convex combination case and compute the explicit asymptotical radius applying Biane and Lehner method. We rewrite
$\mathrm{b}_{\mathrm{c}}=(1-r)u+rq=(1-r)\left(u+\frac{r}{1-r}q\right)=(1-r)(u+q^{\prime})$
@@
where $\varphi=r/(1-r)$ (see [BL99, Equation (4.5)]). Therefore considering $r\leq 1/3$ we have $\rho_{n,\mathrm{c}}={\mathrm{N}}_{n,\mathrm{c}}$. Besides, for all $\epsilon>0$ we have for $n$ large enough ${{\mathrm{N}}}_{n,\mathrm{c}}\leq\rho(p)+\epsilon$ which proves a strictly positive lower bound for the spectral gap for $N$ large enough.
@@
Figure 2: In blue the eigenvalues of P=(1r)M+rQP=(1-r)M+rQ with $Q$ perfect matching and $N=600$. In red the spectral radius of (1r)u+rq(1-r)u+rq.